# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ B.1 Full Results for the Case Studies
Method $\mathrm{Acc}_{Y}$ $\mathrm{F_{1}}(Y)$ $\mathrm{Acc}_{C}$ $\mathrm{F_{1}}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $0.99\pm 0.01$ $0.99\pm 0.01$ $0.65\pm 0.05$ $0.59\pm 0.06$ $0.08\pm 0.07$- -- $0.06\pm 0.01$
CBNM noent $0.99\pm 0.01$ $0.99\pm 0.01$ $0.22\pm 0.01$ $0.08\pm 0.02$ $0.76\pm 0.06$ $0.52\pm 0.04$ $0.52\pm 0.04$ $0.05\pm 0.01$
CBNM 20 $0.99\pm 0.01$ $0.99\pm 0.01$ $0.21\pm 0.01$ $0.07\pm 0.01$ $0.80\pm 0.01$ $0.51\pm 0.01$ $0.52\pm 0.01$ $0.07\pm 0.04$
CBNM 20 noent $0.99\pm 0.01$ $0.99\pm 0.01$ $0.22\pm 0.01$ $0.07\pm 0.01$ $0.80\pm 0.01$ $0.49\pm 0.04$ $0.48\pm 0.04$ $0.04\pm 0.01$
DSL $0.94\pm 0.03$ $0.94\pm 0.03$ $0.20\pm 0.01$ $0.07\pm 0.01$ $0.80\pm 0.01$ $0.02\pm 0.01$ $0.09\pm 0.02$ $0.29\pm 0.16$
bears $0.99\pm 0.01$ $0.99\pm 0.01$ $0.20\pm 0.01$ $0.07\pm 0.01$ $0.78\pm 0.04$ $0.48\pm 0.04$ $0.49\pm 0.04$ $0.03\pm 0.01$
Figure 5:  Fig. 3 with standard deviation over $5$ seeds
Figure 6: Mitigation strategies evaluated for CBNM on Clevr, with standard deviation over $5$ seeds. We evaluated only the contrastive strategy as it performed best. Since supervising all knowledge is infeasible, we observe that for sampled configurations (leading to out-of-distribution settings), the learned knowledge does not generalize.
@@
Figure 7: Confusion matrices of concepts for CBNM on MNIST-Add (first row: before alignment, second row: after alignment) along with the corresponding knowledge: number in cells refer to the predicted sum.
@@
DPL
Refer to caption Refer to caption
DSL
Refer to caption Refer to caption
DPL
Refer to caption Refer to caption
Figure 9: Confusion matrices between ground-truth and predicted concepts. In MNIST-SumParity models tend to favor $\bm{\mathrm{\alpha}}$’s that collapse concepts together, hinting at simplicity bias.
B.2 Number of Reasoning Shortcuts and Joint Reasoning Shortcuts
Here, we report the approximate number of deterministic reasoning shortcuts and joint reasoning shortcuts for MNIST-Add and MNIST-SumParity. We obtained these numbers by extending the count-rss software included in rsbench [Bortolotti et al., 2024] to the enumeration of JRSs, compatibly with Eq. 8 and Eq. 7.
We considered downsized versions of both MNIST-Add and MNIST-SumParity, with input digits restricted to the range $\{0,\ldots,N\}$, and resorted to the approximate model counter approxMC [Chakraborty et al., 2016] for obtaining approximate counts in the hardest cases. All counts assume object independence is in place (i.e., that the two input digits are processed separately), for simplicity. The situation is similar when this is not the case, except that the all counts grow proportionally.
The number of RSs and JRSs for increasing $N$ are reported in Table 12. Specifically, the #RSs column indicates the number of regular RSs obtained by fixing the prior knowledge (that is, assuming $\bm{\mathrm{\beta}}=\bm{\mathrm{\beta}}^{*}$) and matches the count in [Marconato et al., 2023b] for the same tasks. The #JRSs columns refer to the CBM case, where the inference layer is learned from data (i.e., $\bm{\mathrm{\beta}}$ might not match $\bm{\mathrm{\beta}}^{*}$). We report two distinct counts for JRSs: one assumes that $\bm{\mathrm{\beta}}$’s that match on all concepts actively output by $\bm{\mathrm{\alpha}}$ should be considered identical despite possible differences on “inactive” concepts that $\bm{\mathrm{\alpha}}$ predicts as constant (#JRSs non-redundant) or not (# JRSs).
As can be seen from the results, JRSs quickly outnumber regular RSs as $N$ grows, as anticipated in Section 3.1. This suggests that, as expected, learning CBMs with intended semantics by optimizing label likelihood is more challenging when prior knowledge is not given a priori.
@@ B.2 Number of Reasoning Shortcuts and Joint Reasoning Shortcuts
$N$RSs #JRSs #JRSs (non-redundant)
@@ Appendix C Theoretical Material
In this section, we start by summarizing the prerequisite material. We begin by recalling a central Lemma proven in [Marconato et al., 2023b] that allows us to write the condition for likelihood-optimal models in terms of $\bm{\mathrm{\alpha}}$. To ease of comparison, let $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}}):=\bm{\mathrm{f}}(\bm{\mathrm{X}})$ and $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}}):=(\bm{\mathrm{\omega}}\circ\bm{\mathrm{f}})(\bm{\mathrm{X}})$ denote the conditional distributions defined by the CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$, and:
@@
where $\mathsf{KL}$ is the Kullback-Leibler divergence and H is the Shannon entropy. Moreover, under Assumptions 3.1 and 3.2, $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}};\mathsf{K})$ is an optimum of the LHS of Eq. 14 if and only if $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ is an optimum of the RHS. (ii) There exists a bijection between the deterministic concept distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ that, for each $\bm{\mathrm{g}}\in\mathrm{supp}(\bm{\mathrm{G}})$, are constant over the support of $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ apart for zero-measure subspaces $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$ and the deterministic distributions of the form $p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$.
Here, point (i) connects the optima of the likelihood $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}})$ and the optima of $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$. This implies that, under Assumptions 3.1 and 3.2, knowing the optima of $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ informs us about optimal CBMs $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}})$. Notice that, to a single $\bm{\mathrm{\alpha}}(\bm{\mathrm{G}})=p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$ might correspond to multiple choices of $\bm{\mathrm{f}}\in\mathcal{F}$. The choice is unique if, by point (ii), we consider deterministic $p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$ showing that they are into one-to-one correspondence with $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ that are almost constant in the support.
Limiting to those $\bm{\mathrm{\alpha}}$’s in point (ii), we can find a simple way to describe optima for $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$.
@@
Under Assumptions 3.1 and 3.2, for all CBMs $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ where $\bm{\mathrm{f}}:\bm{\mathrm{x}}\mapsto\bm{\mathrm{c}}$ is constant over the support of $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ apart for a zero-measure subspace $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$, for all $\bm{\mathrm{g}}\in\mathcal{G}$, it holds that the optima of the likelihood for $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ are obtained when:
@@
where $\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}):=\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{f}}(\bm{\mathrm{x}})]\in\mathsf{Vert}(\mathcal{A})$ and $\bm{\mathrm{\beta}}(\bm{\mathrm{c}}):=\bm{\mathrm{\omega}}{(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})}\in\mathsf{Vert}(\mathcal{B})$.
@@
Notice that by Lemma C.1, it holds that $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ is a deterministic conditional distribution. Since $\bm{\mathrm{f}}:\bm{\mathrm{x}}\to\bm{\mathrm{c}}$ is constant over all $\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ for a fixed $\bm{\mathrm{g}}\in\mathcal{G}$, it holds that $\bm{\mathrm{f}}(\bm{\mathrm{x}})=\bm{\mathrm{\alpha}}(\bm{\mathrm{g}})$ apart for some $\bm{\mathrm{x}}$ in a measure-zero subspace $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$, therefore we get:
$\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{\omega}}(\bm{\mathrm{f}}(\bm{\mathrm{x}}))]$ $=\bm{\mathrm{\omega}}(\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}))$ (16)
$=(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})$ (17)
where in the last line we substituted $\bm{\mathrm{\omega}}$ with $\bm{\mathrm{\beta}}$ since $\bm{\mathrm{\alpha}}$ is deterministic. For $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$, the necessary and sufficient condition for maximum log-likelihood is:
@@
We start by proving that Definition 3.3 defines an equivalence relation.
@@
Definition 3.3 defines an equivalence relation (denoted as $\sim$) between couples $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}}),(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\in\mathcal{A}\times\mathcal{B}$, where $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})$ if there exist a permutation $\pi:[k]\to[k]$ and $k$ element-wise invertible functions $\psi_{1},\ldots,\psi_{k}$ such that:
@@
To the aim of the proof, it is useful to analyze how $\bm{\mathrm{P}}_{\pi}$ and $\bm{\mathrm{\psi}}$ relate. Consider for any $\bm{\mathrm{c}}\in\mathcal{C}$:
@@
Denoting with $\bm{\mathrm{\psi}}_{\pi}(\bm{\mathrm{c}}):=(\psi_{\pi(1)}(c_{1}),\ldots,\psi_{\pi(k)}(c_{k}))$, we have that:
@@
From this expression, notice it holds for all $\bm{\mathrm{c}}\in\mathcal{C}$:
@@
Symmetry. We have to prove that $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\implies(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\sim(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$. To this end, we start by taking the expression for $\bm{\mathrm{\alpha}}$ in terms of $\bm{\mathrm{\alpha}}^{\prime}$. Consider for any $\bm{\mathrm{g}}\in\mathcal{G}$:
@@
where in the last step we used Eq. 24. By first inverting $\bm{\mathrm{\psi}}_{\pi}$ and then $\bm{\mathrm{P}}_{\pi}$, we obtain that:
@@
showing the reflexivity of $\bm{\mathrm{\alpha}}$. With similar steps, we can show that a similar relation also holds for $\bm{\mathrm{\beta}}^{\prime}$, that is for all $\bm{\mathrm{c}}\in\mathcal{C}$:
@@
for all $\bm{\mathrm{c}}\in\mathcal{C}$. This shows the equivalence relation of Definition 3.3. ∎
We now prove how solutions with intended semantics (Definition 3.3) relate to the optima of the log-likelihood:
@@
If a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ possesses the intended semantics (Definition 3.3), it holds:
@@
We make use of the fact that by Definition 3.3 it holds:
@@
Therefore, combining Eqs. 37 and 38 we get for all $\bm{\mathrm{g}}\in\mathcal{G}$:
@@
Let $s\mathcal{G}:=\bigcup_{i=1}^{k}\{|\mathcal{G}_{i}|\}$ be the set of cardinalities of each concept $G_{i}\subseteq\bm{\mathrm{G}}$ and $ms\mathcal{G}$ denote the multi-set $ms\mathcal{G}:=\{(|\mathcal{G}_{i}|,m(|\mathcal{G}_{i}|),\;i\in[k]\}$, where $m(\bullet)$ denotes the multiplicity of each element of $s\mathcal{G}$. Under Assumptions 3.1 and 3.2, the number of deterministic JRSs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ amounts to:
@@
Since we are considering pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$, we can stick to the fact that $\bm{\mathrm{\alpha}}$ defines a function $\bm{\mathrm{\alpha}}:\mathcal{G}\to\mathcal{C}$ and that, similarly, $\bm{\mathrm{\beta}}$ defines a function $\bm{\mathrm{\beta}}:\mathcal{C}\to\mathcal{Y}$. In this case, $\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}}:\mathcal{G}\to\mathcal{Y}$ is a map from ground-truth concepts to labels. We start by Definition 3.3 and consider the pairs that attain maximum likelihood by Lemma C.2:
@@
where, since $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ and is deterministic, we can substitute $\bm{\mathrm{\omega}}$ with $\bm{\mathrm{\beta}}$. Since both $\mathsf{Vert}(\mathcal{A})$ and $\mathsf{Vert}(\mathcal{B})$ are countable, we can convert this to a total count of pairs that attain maximum likelihood:

### v2

@@ B.1 Additional Results for Q1
Method $\mathrm{Acc}_{Y}$ $F_{1}(Y)$ $\mathrm{Acc}_{C}$ $F_{1}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $9\pm 0.01 0.99±0$$9\pm 0.01$ $0.65\pm 0.05$ $59\pm 0.06 0.0$$8\pm 0.07 - - 0.0$$06\pm 0.01$
[…]
CBNM 20 $9\pm 0.01 0.99±0$$9\pm 0.01$ $21\pm 0.01 0.0$$07\pm 0.01$ $0.80\pm 0.01$ $0.51\pm 0.01$ $52\pm 0.01 0.0$$07\pm 0.04$
[…]
DSL $0.94\pm 0.03$ $0.94\pm 0.03$ $20\pm 0.01 0.0$$07\pm 0.01$ $80\pm 0.01 0.0$$2\pm 0.01 0.0$$09\pm 0.02$ $0.29\pm 0.16$
[…]
B.2 Additional Results for Q1: The BDD-OIA Task
Experimental setup. BDD-OIA is a challenging real-world learning task with $21$ binary concepts and, unlike in Clevr, these cannot be decomposed into unary predicates. Moreover, each concept cannot be separately assigned to individual actions as done for DPL in [22], as this would introduce a negative bias in the learned solutions, effectively forcing the model to use a fixed concept budget per action without interference. Due to scalability constraints, we compared only DPL and CBNM, as implementing DSL or DPL would require allocating at least four tensors of size $2^{21}$, which exceeds our available computational resources.
Results. The results for DPL and CBNM are reported in Table˜14. Both competitors perform sub-optimally on BDD-OIA, as expected given the challenging nature of this task: DPL scores around $70\%$ $F_{1}(Y)$ while CBNM only $62\%$. Perhaps surprising, CBNM does not collapse concepts together and tends to use most of the concept bottleneck across the test set ($\mathsf{Cls}(C)\approx 0.06$). Despite the lack of collapse, both architectures are far from attaining the intended semantics, with $F_{1}(C)$ scores of $44\%$ for DPL and $43\%$ for CBNM. Additionally, CBNM deviates from the ground-truth knowledge, achieving a low $F_{1}(\bm{\mathrm{\beta}})$ of around $42\%$. The low collapse for CBNM may be due to the trained model being far from optimal. We hypothesize that adding a sparsity penalty to the linear layer of CBNM would substantially increase concept collapse, but we leave a detailed examination to future work.
Table 14: Results for BDD-OIA.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.69\pm 0.03$ $0.44\pm 0.01$ $0.88\pm 0.01$ -
[…]
B.3 Additional Results for Q3
Here we show the effect of traditional mitigation strategies on MNIST-SumParity and Clevr as concept and knowledge supervision increase. As discussed in Table˜5, traditional strategies have limited impact, with the exception of concept supervision.
Figure 5: Fig.˜3 with standard deviation over 55 seeds
Figure 6: Effect of standard mitigation strategies evaluated for CBNM on Clevr, with standard deviation over 55 seeds. We evaluated only the contrastive strategy as it performed best. Since supervising all knowledge is infeasible, we observe that for sampled configurations (leading to out-of-distribution settings), the learned knowledge does not generalize.
B.4 Concept Confusion Matrices and Learned Inference Layers
Here, we present the concept and knowledge confusion matrices for different datasets and models to show examples of joint reasoning shortcut solutions that they learn.
@@ B.4 Concept Confusion Matrices and Learned Inference Layers
Figure 7: Concept confusion matrices for CBNM on MNIST-Add and the corresponding learned inference layer. Left two matrices: concepts and knowledge before alignment. Right two matrices: same but after alignment. The learned inference layer is visualized as a colored matrix where the numbers in the cells indicate the model’s predictions.
@@
CBNM DPL DSL DPL
MNIST-Add
Refer to caption Refer to caption Refer to caption Refer to caption
MNIST-SumParity
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Concept confusion matrices for MNIST-SumParity. All models tend to favor $\bm{\mathrm{\alpha}}$’s that collapse concepts together, hinting at simplicity bias.
@@
Figure 10: Concept confusion matrices for SENN on MNIST-Add and MNIST-SumParity. SENN provide local explanations for individual predictions, as their inference layer depends on the input and does not yield a global, interpretable rule set. No simplicity bias is evident in these two confusion matrices; we hypothesize that this stems from the models’ flexibility in modeling local rules and from the reconstruction penalty applied to the concepts during training.
B.5 Number of Reasoning Shortcuts and Joint Reasoning Shortcuts
Here, we report the approximate number of deterministic reasoning shortcuts and joint reasoning shortcuts for MNIST-Add and MNIST-SumParity. We obtained these numbers by extending the count-rss software included in rsbench [22] to the enumeration of JRSs, compatibly with Eq.˜7 and Eq.˜6.
We considered downsized versions of both MNIST-Add and MNIST-SumParity, with input digits restricted to the range $\{0,\ldots,N\}$, and resorted to the approximate model counter approxMC [106] for obtaining approximate counts in the hardest cases. All counts assume object independence is in place (i.e., that the two input digits are processed separately), for simplicity. The situation is similar when this is not the case, except that the all counts grow proportionally.
The number of RSs and JRSs for increasing $N$ are reported in Table˜15. Specifically, the #RSs column indicates the number of regular RSs obtained by fixing the prior knowledge (that is, assuming $\bm{\mathrm{\beta}}=\bm{\mathrm{\beta}}^{*}$) and matches the count in [19] for the same tasks. The #JRSs columns refer to the CBM case, where the inference layer is learned from data (i.e., $\bm{\mathrm{\beta}}$ might not match $\bm{\mathrm{\beta}}^{*}$). We report two distinct counts for JRSs: one assumes that $\bm{\mathrm{\beta}}$’s that match on all concepts actively output by $\bm{\mathrm{\alpha}}$ should be considered identical despite possible differences on “inactive” concepts that $\bm{\mathrm{\alpha}}$ predicts as constant (#JRSs non-redundant) or not (# JRSs).
As can be seen from the results, JRSs quickly outnumber regular RSs as $N$ grows, as anticipated in Section˜3.1. This suggests that, as expected, learning CBMs with intended semantics by optimizing label likelihood is more challenging when prior knowledge is not given a priori.
@@ B.5 Number of Reasoning Shortcuts and Joint Reasoning Shortcuts
NN #RSs #JRSs #JRSs (non-redundant)
@@ Appendix C Theoretical Material
In this section, we start by summarizing the prerequisite material. We begin by recalling a central Lemma proven in [19] that allows us to write the condition for likelihood-optimal models in terms of $\bm{\mathrm{\alpha}}$. For ease of comparison, let $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}}):=\bm{\mathrm{f}}(\bm{\mathrm{X}})$ and $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}}):=(\bm{\mathrm{\omega}}\circ\bm{\mathrm{f}})(\bm{\mathrm{X}})$ denote the conditional distributions defined by the CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$, and:
@@
where $\mathsf{KL}$ is the Kullback-Leibler divergence and H is the Shannon entropy. Moreover, under Assumptions˜3.1 and 3.2, $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}};\mathsf{K})$ is an optimum of the LHS of Eq.˜17 if and only if $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ is an optimum of the RHS. (ii) There exists a bijection between the deterministic concept distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ that, for each $\bm{\mathrm{g}}\in\mathrm{supp}(\bm{\mathrm{G}})$, are constant over the support of $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ apart for zero-measure subspaces $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$ and the deterministic distributions of the form $p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$.
Point (i) connects the optima of the likelihood $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}})$ and the optima of $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$. This implies that, under Assumptions˜3.1 and 3.2, knowing the optima of $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ informs us about optimal CBMs $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{X}})$. Notice that a single $\bm{\mathrm{\alpha}}(\bm{\mathrm{G}})=p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$ might correspond to multiple choices of $\bm{\mathrm{f}}\in\mathcal{F}$. However, by (ii), if we restrict ourselves to deterministic distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{G}})$, the correspondence with distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ that are almost constant in the support becomes one-to-one.
Next, we describe which distributions $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ are both deterministic – and thus comply with (ii) – and optimal.
@@
Under Assumptions˜3.1 and 3.2, for all CBMs $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ where, for all $\bm{\mathrm{g}}\in\mathcal{G}$, $\bm{\mathrm{f}}:\bm{\mathrm{x}}\mapsto\bm{\mathrm{c}}$ is constant over the support of $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ apart for a zero-measure subspace $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$, it holds that the optima of the likelihood for $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$ are obtained when:
@@
Here, $\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}):=\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{f}}(\bm{\mathrm{x}})]\in\mathsf{Vert}(\mathcal{A})$ and $\bm{\mathrm{\beta}}(\bm{\mathrm{c}}):=\bm{\mathrm{\omega}}{(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})}\in\mathsf{Vert}(\mathcal{B})$.
@@
Notice that by Lemma˜C.1, it holds that $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ is a deterministic conditional distribution. Since $\bm{\mathrm{f}}:\bm{\mathrm{x}}\to\bm{\mathrm{c}}$ is constant over all $\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$ for a fixed $\bm{\mathrm{g}}\in\mathcal{G}$, it holds that $\bm{\mathrm{f}}(\bm{\mathrm{x}})=\bm{\mathrm{\alpha}}(\bm{\mathrm{g}})$ apart for some $\bm{\mathrm{x}}$ in a measure-zero subspace $\mathcal{X}^{0}\subset\mathsf{supp}{(p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}}))}$. Therefore we get:
$\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{\omega}}(\bm{\mathrm{f}}(\bm{\mathrm{x}}))]=\bm{\mathrm{\omega}}(\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}))=(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})$ (19)
In the last line, we substitute $\bm{\mathrm{\omega}}$ with $\bm{\mathrm{\beta}}$ since $\bm{\mathrm{\alpha}}$ is deterministic. For $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}})$, the necessary and sufficient condition for maximum log-likelihood is:
@@
We start by proving that Definition˜3.3 defines an equivalence relation.
@@
Definition˜3.3 defines an equivalence relation $\sim$ between pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$, $(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\in\mathcal{A}\times\mathcal{B}$, as follows: $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})$ iff there exist a permutation $\pi:[k]\to[k]$ and $k$ element-wise invertible functions $\psi_{1},\ldots,\psi_{k}$ such that:
@@
It is useful to analyze how $\bm{\mathrm{P}}_{\pi}$ and $\bm{\mathrm{\psi}}$ are related. Let for any $\bm{\mathrm{c}}\in\mathcal{C}$:
@@
Using the shorthand $\bm{\mathrm{\psi}}_{\pi}(\bm{\mathrm{c}}):=(\psi_{\pi(1)}(c_{1}),\ldots,\psi_{\pi(k)}(c_{k}))$, we have that:
@@
From this expression, notice that for all $\bm{\mathrm{c}}\in\mathcal{C}$ it holds that:
@@
Symmetry. We have to prove that $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\implies(\bm{\mathrm{\alpha}}^{\prime},\bm{\mathrm{\beta}}^{\prime})\sim(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$. Since $(\alpha,\beta)\sim(\alpha^{\prime},\beta^{\prime})$, we can write $\alpha$ in terms of $\alpha^{\prime}$ as follows:
@@
where in the last step we used Eq.˜26. By first inverting $\bm{\mathrm{\psi}}_{\pi}$ and then $\bm{\mathrm{P}}_{\pi}$, we obtain that:
@@
The inverses exist by construction/definition. Showing the symmetry of $\bm{\mathrm{\alpha}}$. With similar steps, we can show that a similar relation also holds for $\bm{\mathrm{\beta}}^{\prime}$, that is for all $\bm{\mathrm{c}}\in\mathcal{C}$:
@@
for all $\bm{\mathrm{c}}\in\mathcal{C}$. This shows the equivalence relation of Definition˜3.3. ∎
We now prove how solutions with intended semantics (Definition˜3.3) relate to the optima of the log-likelihood:
@@
If a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ possesses the intended semantics (Definition˜3.3), it holds that:
@@
By Definition˜3.3, we can write:
@@
Combining Eqs.˜39 and 40, we get, for all $\bm{\mathrm{g}}\in\mathcal{G}$, that:
@@
Let $s\mathcal{G}:=\bigcup_{i=1}^{k}\{|\mathcal{G}_{i}|\}$ be the set of cardinalities of each concept $G_{i}\subseteq\bm{\mathrm{G}}$ and $ms\mathcal{G}$ denote the multi-set $ms\mathcal{G}:=\{(|\mathcal{G}_{i}|,m(|\mathcal{G}_{i}|)),\;i\in[k]\}$, where $m(\bullet)$ denotes the multiplicity of each element of $s\mathcal{G}$. Under Assumptions˜3.1 and 3.2, the number of deterministic JRSs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ amounts to:
@@
Since we are considering pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$, $\bm{\mathrm{\alpha}}$ defines a (deterministic) function $\bm{\mathrm{\alpha}}:\mathcal{G}\to\mathcal{C}$ and that, similarly, $\bm{\mathrm{\beta}}$ defines a (deterministic) function $\bm{\mathrm{\beta}}:\mathcal{C}\to\mathcal{Y}$. In this case, $\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}}:\mathcal{G}\to\mathcal{Y}$ is a map from ground-truth concepts to labels. We start by Definition˜3.3 and consider the pairs that attain maximum likelihood by Lemma˜C.2:
@@
Since $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ is deterministic, we can replace $\bm{\mathrm{\omega}}$ with $\bm{\mathrm{\beta}}$. Since both $\mathsf{Vert}(\mathcal{A})$ and $\mathsf{Vert}(\mathcal{B})$ are countable, we can count the number of pairs that attain maximum likelihood as follows: